1. You want to buy a $24,000 car. The company is offering a 5% interest rate for 48 months (4 years). What will your monthly payments be?

2. You have $3,500 on a credit card that charges a 19% interest rate. If you want to pay off the credit card in 5 years, how much will you need to pay each month (assuming you don't charge anything new to the card)?

$ each month
Answer #1

1.
Using financial calculator
PV=-24000
N=48
I/Y=5%/12
FV=0
CPT PMT=552.70

2.
Using financial calculator
PV=-3500
N=12*5
I/Y=19%/12
FV=0
CPT PMT=90.79